OpenStudy (anonymous):

Tell whether the sequence is arithmetic. If it is identify the common difference 30. -15,-5,5,15….

OpenStudy (anonymous):

25,

OpenStudy (anonymous):

is there a common difference? -15 plus what =-5

OpenStudy (anonymous):

-15+10=-5, -5+10=+5, +5+10=15, then +15+10=25, you have to add 10 to each obtained value in each sequences. Does it help?